A -> B -> C

At time t=0 element A has 50,000,000 particles and element B has 10,000,000 particles. Element A has a half-life of four months and element B has a half-life of seven months. How many particles of A and B will there be one year later?

Secular Equilibrium

At t=0 there are 10,000,000 atoms of element A. It has a half-life of five days. There are no atoms of element B or C. Element B has a one-day half-life, C a one-hour half-life. After two days, approximately how many atoms of each element are there? Note that the math for this problem should be a great deal easier than the math for the first problem.
Answer #1

We know that

N = No e-kt where k is the decay constant given by k = 0.693/t1/2

For A,

N = 50,000,000(0.128) = 6443,185

Similarly For B,

N = 10,000,000(0.334) = 3344727